Assume the car can be purchased for 0% down for 60 months (in lieu of rebate). A BMW that has a sticker price of $62,430 with factory and dealer rebates of $6,000 (a) Find the monthly payment if financed for 60 months at 0% APR. (Round your answer to the nearest cent.) $ 46.64 Incorrect: Your answer is incorrect. (b) Find the monthly payment if financed at 2.5% add-on interest for 60 months. (Round your answer to the nearest cent.) $ 408.31 Incorrect: Your answer is incorrect. (c) Use the APR approximation formula to find the APR for part (b). (Round your answer to one decimal place.) 50.747 Incorrect: Your answer is incorrect. % (d) State whether the 0% APR or the 2.5% add-on rate should be preferred. 0% APR 2.5% add-on rate Correct: Your answer is correct.
Answer #1

Part A

Monthly payment = 62430/60 = $1040.50

Part B

discount price = 62430-6000 = 56430

add-on-interest = 56430*2.5%*60/12 = 7053.75

total value = 56430+7053.75 = 63483.75

monthly payment = 63483.75/60 = $1058.06

Part C

APR = 2Nr/(N+1) = (2*60*2.5%)/(60+1) = 4.9%

Part D

0% APR should be preferred

Because monthly payment is lower in that case.